Televisions sets; What does it mean to refer to a 20 inch TV set or a 25 inch TV set? Such units refer to the diagonal of the screen. A 25 inch TV set also has a width of 20 inches. What is its height?
Answer by checkley77(12844)

A^2+B^2=C^2
A^2+20^2=25^2
A^2+400=625
A^2=625-400
A^2=225
A=SQRT225
A=15 INCHES IS THE HEIGHT.